A true-false test consisting of 20 questions can be marked in 1,048,576 different ways. In how many ways can each question be marked true or false so that
a) 7 are right and 13 are wrong
b) at least 17 are right?
Thanks!

Answer by stanbon(75887)

A true-false test consisting of 20 questions can be marked in 1,048,576 different ways. In how many ways can each question be marked true or false so that
---
Binomial Problem with n = 20 and p = 1/2
------
a) 7 are right and 13 are wrong
P(x = 7) = 20C7(1/2)^20 = 77520/1048576 =0.0739
------------------
b) at least 17 are right?
P(17<= x <=20) = 1 - binomcdf(20,0.5,16) = 0.001288
----------------
Cheers,
Stan H.